# Bayesian dictionary learning estimation of cell membrane permeability from surface pH data

### v1

@@ Abstract
Gas transport across cell membrane is a very important process in biochemistry which is essential for many crucial tasks, including cell respiration pH regulation in the cell. In the late 1990’s, the suggestion that gasses are transported via preferred gas channels embedded into the cell membrane challenged the century old Overton’s theory that gases pass through the lipid cell membrane by diffusing across the concentration gradient. Since experimental evidence alone does not provide enough evidence to favor one of the proposed mechanisms, mathematical models have been introduced to provide a context for the interpretation of laboratory measurement. Following up on previous work where the membrane permeability was estimated using particle filter, in this article we propose an algorithm based on dictionary learning for estimating cell membrane permeability. Computed examples illustrate that the novel approach, which can be applied when the properties of the membrane do not change in the course of the data collection process, is computationally much more efficient than particle filter.
@@ 1 Introduction
Gas transport across cell membrane has been studied extensively because of its essential role in many cell functions, including cell respiration and pH regulation. In the literature on acid-base physiology, the regulation of the exchange of carbon dioxide (${\rm CO}_{2}$) between the cell and its surrounding environment has been a recurrent topic, because the shuttle mechanism effectuating the passage, shown schematically in Figure 1, is one of the ways for the cell to regulate the pH near the membrane surface.
@@
The mechanism of gas exchange through cell membrane and the properties of membrane permeability have been investigated since the 1890’s, starting with Overton’s experiments and his theory [26], based on the hypothesis that the gases diffuse through the lipidic cell membrane driven by the concentration gradient across the membrane. In the late 1990s a new hypothesis emerged, suggesting with mounting experimental evidence that, in addition to diffusion, specialized gas channels, namely aquaporins (AQPs) and Rhesus proteins (Rh), might play a significant role in the exchange of gasses between cytoplasm and the outside environment [5, 21, 25, 15]. These gas channels, which are protein complexes embedded in membranes of various kinds of cells, were postulated to allow selectively the exchange of certain gases across the cell membrane [22, 17, 18].
@@
To validate the gas channel hypothesis, several experimental studies have been published on the passage of ${\rm CO}_{2}$ across the membrane of the oocytes of Xenopus laevis [13, 25, 22], a cell large enough to make it possible to inject it with heterologous RNA so as to have it express different kinds of AQPs and Rh proteins on its membrane, thus allowing to investigate the effects of the AQPs (or Rh) on the cell membrane permeability to gases [15]. The mathematical model proposed in [27] describes the standard experiment with the oocytes of X. laevis, in which the cell is immersed in an aqueous solution with high concentration of ${\rm CO}_{2}$, and the pH value directly above the surface is measured by means of a micro-sensor. The mathematical model, which includes a biochemical description of the shuttle mechanism relating the passage of ${\rm CO}_{2}$ and changes in pH, provides a context for the interpretation of the experimental findings. In this model, it is assumed that the only chemical species that passes through the X. laevis cell membrane is ${\rm CO}_{2}$, while all other chemical species that are present in the solution cannot cross the membrane which is de facto impermeable to them. This hypothesis is supported by earlier studies, see, e.g., [16, 14].
@@
The standard experiment comprises the measurement of the surface pH on the cell membrane by a pH-sensitive electrode placed against the membrane. At the beginning of the experiment, the data show a steep rise of the surface pH caused by the association of protons and bicarbonate into carbonic acid that breaks into water and ${\rm CO}_{2}$ replacing the carbon dioxide passing through the membrane into the cell. As the experiment proceeds, the ${\rm CO}_{2}$ transport slows down as the concentration inside the cell rises towards that of the outside, and this, together with the diffusion from the exterior domain drives the system to a new equilibrium. The virtual proton transport into the cell is enhanced by the presence of carbonic anhydrases (CA), a family of enzymes that increase the dissociation of the ${\rm H}_{2}{\rm CO}_{3}$ right outside the cell membrane, while promoting the reverse reaction right under the membrane surface [5, 21, 24, 23], as illustrated schematically in Figure 1. The concentration and spatial distribution of CA varies for different kind of cells, and it is known to be significantly high, for example, in erythrocytes.
@@
The computational complexity of the detailed forward model [7] makes it unsuitable for solving the inverse problem of estimating the permeability of the membrane from the pH data, thus a reduced approximate model still accounting for the effect of the sensor was proposed in [2], and subsequently used for the estimation of the membrane permeability as the evolution step of a Bayesian particle filter algorithm in [3].
@@
In recent years, there has been a growing interest in using dictionary learning methods for solving inverse problems when the data are not sufficient or the forward model is not amenable for classical optimization-based approaches. Following this idea, here we propose a new alternative algorithm for solving the inverse problems of estimating the cell membrane properties based on a dictionary learning algorithm introduced in [4]. The main advantage of the new approach is that it is computationally much more efficient than the particle filter algorithm, estimating the desired parameters to a similar accuracy as the particle filter at a small fraction of the computing time.
@@ 2 Experimental framework and mathematical model
The experiment start with a homogeneous solution containing carbon dioxide (${\rm CO}_{2}$), bicarbonate (${\rm HCO}_{3}^{-}$), carbonic acid (${\rm H}_{2}{\rm CO}_{3}$), hydrogen ion (${\rm H}^{+}$), a non-specified pH buffer other than bicarbonate (${\rm A}^{-}$), and its the relative acid (${\rm HA}$) in equilibrium with known initial concentrations. In the standard experiment, the X. leavis oocyte is immersed in this solution, and the micro-electrode measuring the membrane surface pH is already in place. The interior of the cell is assumed to consist of a solution of the same species as the exterior, but with different known equilibrium concentrations.
@@
In [7], where the micro-environment under the tip of the sensor was modeled using a 3D finite element method approach, a substantial computational effort was required for each simulation, highlighting the need for a lower complexity model for the solution of the inverse problem of membrane permeability. Following [2], here we resort to a lumped compartment model for the perturbation caused by the sensor tip near the cell membrane.
@@
where $u_{i}^{\pm}=u_{i}^{\pm}(x,t)$ is the concentration of the $i$th species at time $t$ and point $x$, the superscript $\pm$ indicates whether the concentration refers to that inside (-) or the outside (+) the cell membrane, $\varphi^{\pm}_{nj}=\varphi^{\pm}_{nj}(x,t)$ is the reaction flux of the $j$th reaction and, finally, $S_{nj}$ are the entries of the stoichiometric matrix $S$. More specifically, the numbering of the concentrations is $u_{1}=[{\rm CO}_{2}]$, $u_{2}=[{\rm H_{2}CO}_{3}]$, $u_{3}=[{\rm HCO}_{3}^{-}]$, $u_{4}=[{\rm H}^{+}]$, $u_{5}=[{\rm HA}]$, and $u_{6}=[{\rm A}^{-}]$. The diffusion coefficients $\kappa_{n}^{\pm}$, assumed to be known, are set to the values reported in the literature listed in Table 3.
@@
Substance Inside [$(\mu\,{\rm m})^{2}/s$] Outside [$(\mu\,{\rm m})^{2}/s$]
${\rm CO_{2}}$ 1.71×$\times 10^{3}$ 1.71×$\times 10^{3}$
${\rm H_{2}CO_{3}}$ 1.11×$\times 10^{3}$ 1.11×$\times 10^{3}$
${\rm HCO_{3}}^{-}$ 1.11×$\times 10^{3}$ 1.11×$\times 10^{3}$
${\rm H}^{+}$ 8.69×$\times 10^{3}$ 8.69×$\times 10^{3}$
${\rm HA}$ 1.56×$\times 10^{3}$ 1.56×$\times 10^{3}$
${\rm A}^{-}$ 1.56×$\times 10^{3}$ 1.56×$\times 10^{3}$
@@
At the beginning of the experiment, the systems inside and outside the membrane are in equilibrium with concentrations constants throughout the exterior domain and within the cell, however, the constant concentrations are not equal across the membrane boundary. The initial values of the concentrations are listed in Table 3. The assumption that solution is well-mixed solution and the concentrations of the species far from the cell are well-known justifies assigning Dirichlet boundary conditions at an artificial outer boundary of the spherical computational domain with radius $R_{\infty}$,
@@
Reaction $k_{\ell}$ $k_{-\ell}$ $K=k_{\ell}/k_{-\ell}$
${\displaystyle{\rm CO}_{2}+{\rm H}_{2}{\rm O}{\mathop{\rightleftharpoons}^{k_{1}}_{k_{-1}}}{\rm H}_{2}{\rm CO}_{3}}$ 0.0302 [1/s] 10.9631 [1/s] 2.7547×$\times 10^{-4}$
${\displaystyle{\rm H}_{2}{\rm CO}_{3}{\mathop{\rightleftharpoons}^{k_{2}}_{k_{-2}}}{\rm HCO}_{3}^{-}+{\rm H}^{+}}$ $\varepsilon=10^{-9}$ [1/s] $\varepsilon/K_{2}$ $K_{2}=0.2407$ [mM]
${\displaystyle{\rm HA}_{\rm in}{\mathop{\rightleftharpoons}^{k_{3}}_{k_{-{3}}}}{\rm A}^{-}_{\rm in}+{\rm H}^{+}_{\rm in}}$ $\varepsilon^{\prime}=10^{-6}$ [1/s] $\varepsilon^{\prime}/K_{\rm HA}$ $K_{\rm HA}=7.9433\times 10^{-5}$ [mM]
${\displaystyle{\rm HA}_{\rm out}{\mathop{\rightleftharpoons}^{k_{3}}_{k_{-{3}}}}{\rm A}^{-}_{\rm out}+{\rm H}^{+}_{\rm out}}$ $\varepsilon^{\prime}=10^{-6}$ [1/s] $\varepsilon^{\prime}/K_{\rm HA}$ $K_{\rm HA}=3.1623\times 10^{-5}$ [mM]
Table 4: Reaction rates. Note that $\varepsilon$ and $\varepsilon^{\prime}$, whose precise values are not well known, represent the fast rates of the reactions for which no better guess is available. Furthermore, in the simulations we use $Ak_{\pm 1}$, as those reactions are enhanced by the (multiplicative) acceleration factor $A$.
@@
[…]
@@ 3 Bayesian dictionary learning
In data science, dictionary learning and dictionary matching refer to a class of algorithms aiming at explaining a given datum in terms of a collection of precomputed or previously measured and labeled entries, referred to as the underlying dictionary. In applications where the new datum is a vector $b$ of measured response of a physical system, dictionary learning can be employed to explain it as a a linear combination of a collection of possible responses. Assuming that the known responses are stored as vectors $d_{1},\dots,d_{p}$ of the same dimension as $b$, often referred to as atoms, they can be organized as the columns of a matrix dictionary ${\mathsf{D}}$, and the problem can be formulated as finding a solution of the linear system
@@
In inverse problem applications, the atoms may represent precomputed outputs corresponding to a model that depends on model parameters that one is primarily interested in, the model parameter values serving as labels. Identifying one or few atoms that are able to explain the datum allows an identification of the model parameters, thus providing a solution to the inverse problem. The approach is particularly useful for problems in which the forward model is not amenable for classical methods, see, e.g., [20, 1] for examples.
@@
Assuming that the dictionary atoms are partitioned into $K$ clusters, refereed to as sub-dictionaries, following [4], we can permute the columns of ${\mathsf{D}}$ so that columns corresponding to atoms in the same cluster are adjacent, yielding the block structure
@@
In this manner the dictionary learning problem can be solved by first identifying the sub-dictionaries that may be pertinent, and subsequently reformulating (20) with a smaller dictionary matrix that does not include the blocks corresponding to irrelevant sub-dictionaries. It has been shown in [4] that this approach can substantially reduce the computational complexity, especially when the dictionary comprises a large number of atoms. If there is no obvious way of partitioning the atoms in subdictionaries, a block partitioning of the columns of the dictionary matrix ${\mathsf{D}}$ can be obtained via unsupervised clustering using, e.g., $k$-means or $k$-medoids algorithm.
In data science applications, it is customary to reduce the size of large data matrices by assessing their effective dimensionality, typically by computing a singular value decomposition, and then addressing the data in terms of its principal components corresponding to the largest singular values, a technique known as Principal Component Analysis (PCA). Once the sub-dictionaries have been identified, e.g., by clustering, each one can be compressed and approximated by just a few feature vectors. More specifically, if ${\mathsf{D}}^{(i)}\in{\mathbb{R}}^{m\times p_{i}}$, we write a low-rank approximation
@@
that is, every atom in the sub-dictionary ${\mathsf{D}}^{(i)}$ is approximated by a linear combination of the columns of $W^{(i)}$, referred to as the $i$th code book. The rank $k_{i}$ represents the effective dimensionality of the subdictionary, and $W^{(i)}$ represents a reduced sub-dictionary because it summarizes ${\mathsf{D}}^{(i)}$ in terms of much fewer feature vectors. The low-rank approximation in PCA is based on the singular value decomposition (SVD) of the matrix, however, if positivity constraints for the feature vectors and/or the coefficients are necessary, non-negative matrix factorization (NMF) algorithms may be preferable [19]. While for a given approximation rank $k_{i}$, the approximation properties of NMF measured in the Frobenius norm are inferior to those of PCA, the non-negativity of the feature vectors and of the weights facilitate the interpretation of the factorization.
@@
the inequalities being understood component-wise. The columns of the matrix ${\mathsf{W}}^{(i)}$, sometimes referred to as feature vectors, can be interpreted as a summary of the atoms of ${\mathsf{D}}^{(i)}$, thus ${\mathsf{W}}^{(i)}$ can be thought of as a compressed version of ${\mathsf{D}}^{(i)}$.
@@
$D^{(i)}=W^{(i)}H^{(i)}+{\mathsf{E}}^{(i)}\,.$ (25)
In the proposed algorithm, the dictionary compression error is interpreted as an unknown noise term with an underlying probability distribution. Following the Bayesian paradigm where all unknowns, including modeling errors, are interpreted as random variables, we assume that the compression error vectors are realizations of a random variable $E^{(i)}$, and estimate the underlying probability distribution by sampling realizations of this random variable. Therefore, for each subdictionary, we generate a sample of dictionary entries $\{d_{j}^{(i)}\}_{j=1}^{K_{i}}$, find their approximate representation in terms of the feature vectors, and compute the approximation errors,
@@
[…]
@@
The final step of the process entails going from the solution vector $\widehat{x}$ to the parameter vector $\xi$ associated with the measured data $b$. If $\widehat{x}$ contains only a single non-zero component, the parameter identification is straightforward. When the sparse vector $x$ contains more than one non-zero entries, a winner-takes-all strategy of electing the parameter vector corresponding to the best matching dictionary atom is a feasible approach. In the computed examples of this paper, we use a heuristic interpolation scheme. Consider the mapping from the parameter space to the data space, $f:\xi\mapsto d$. We denote by $\xi_{j}^{(\widehat{i})}$ the parameter vectors corresponding to the best matching subdictionary, $1\leq j\leq p_{\,\widehat{i}}$, and by $d_{j}^{(\widehat{i})}$ the corresponding computed outputs. Assuming for a moment that $f$ is approximately linear, we have
$b\approx\sum_{j=1}^{p_{\,\widehat{i}}}\widehat{x}_{j}d^{(\widehat{i})}_{j}=\sum_{j=1}^{p_{\,\widehat{i}}}\widehat{x}_{j}f(\xi_{j}^{(\widehat{i})})\approx f\left(\sum_{j=1}^{p_{\,\widehat{i}}}\widehat{x}_{j}\xi_{j}^{(\widehat{i})}\right),$
leading to the parameter identification
$\widehat{\xi}=\sum_{j=1}^{p_{\,\widehat{i}}}\widehat{x}_{j}\xi_{j}^{(\widehat{i})}.$ (32)
The linearity assumption is in general may not be valid, and the viability of such interpolation scheme should be considered case-by-case.

### v2

@@ Abstract
Gas transport across cell membrane is a very important process in cell biology, as it is essential for many crucial tasks, including cell respiration and pH regulation. Since it is difficult to observe directly how gasses cross cell membranes, mathematical models have played and continue to play an important role in providing a context for the experimental findings depending on gas exchange, and to test different hypotheses about the gas passage mechanism. The first family of gas transport models has been able to match qualitatively the experimental data, and a second generation of computational models has reproduced and explained the dynamic range of the data. These models depend on parameters that cannot be measured experimentally, but can be estimated from the available data and some a priori knowledge about the problem based on first principles of cell physiology. The membrane permeability is arguably one of the most important parameters, and can also be used also to shed light on the validity of different models of the transport mechanisms. In this contribution we propose a new, computationally efficient approach to the cell permeability estimation from measurements of the pH on the cell membrane over the course of a lab experiment. The methodology is inspired by a technique used in data science applications known as dictionary learning.
Relevance to life sciences. In the late 1990’s, the suggestion that gasses are transported via preferred gas channels embedded into the cell membrane challenged the century-old Overton’s theory that gases pass through the lipid cell membrane by diffusing across it driven by the concentration gradient. Since the laboratory experiments alone do not provide enough evidence to favor one or another of the proposed mechanisms, mathematical models have been introduced to provide a context for the interpretation of the measurements. In this article we address the problem of estimating intrinsic properties of the cell membrane that could help resolve the question about the role of aquaporins and other membrane-embedded proteins in gas transport across cell membrane from typical pH measurements on the membrane of the oocytes of Xenopus laevis.
Mathematical content: Following up on a previous work where the membrane permeability was estimated using particle filter techniques, in this article we propose an algorithm based on dictionary learning for estimating cell membrane permeability and other pertinent model parameters. Dictionary learning methods have proven to provide versatile tools to solve inverse problems in which the insufficiency of the available data or the complexity of the forward model makes the use of optimization-based inverse solvers impractical. Computed examples illustrate that the novel approach, which can be applied when the properties of the membrane do not change in the course of the data collection process, is computationally more efficient than the particle filter approach.
@@ 1 Introduction
Gas transport across the cell membrane has been studied extensively because of its essential role in many cell functions, including cell respiration and pH regulation. In the literature on acid-base physiology, the regulation of the exchange of carbon dioxide (${\rm CO}_{2}$) between the cell and its surrounding environment has been a recurrent topic, because the shuttle mechanism effectuating the passage, shown schematically in Figure 1, is one of the ways for the cell to regulate the pH near the membrane surface.
@@
The mechanism of gas exchange through the cell membrane and the properties of membrane permeability have been investigated since the 1890’s, starting with Overton’s experiments and his theory [26], based on the hypothesis that the gases diffuse through the lipidic cell membrane driven by the concentration gradient across the membrane. In the late 1990s a new hypothesis emerged, suggesting with mounting experimental evidence that, in addition to diffusion, specialized gas channels, namely aquaporins (AQPs) and Rhesus proteins (Rh), might play a significant role in the exchange of gasses between cytoplasm and the outside environment [5, 21, 25, 15]. These gas channels, which are protein complexes embedded in membranes of various kinds of cells, were postulated to allow selectively the exchange of certain gases across the cell membrane [22, 17, 18].
@@
To validate the gas channel hypothesis, several experimental studies have been published on the passage of ${\rm CO}_{2}$ across the membrane of the oocytes of Xenopus laevis [13, 25, 22], a cell large enough to make it possible to inject it with heterologous RNA so as to have it express different kinds of AQPs and Rh proteins on its membrane, thus allowing to investigate the effects of the AQPs (or Rh) on the cell membrane permeability to gases [15]. The mathematical model proposed in [27] describes the standard experiment with the oocytes of X. laevis, in which the cell is immersed in an aqueous solution with high concentration of ${\rm CO}_{2}$, and the pH value directly above the surface is measured by means of a micro-sensor. The mathematical model, which includes a biochemical description of the shuttle mechanism relating the passage of ${\rm CO}_{2}$ and changes in pH, provides a context for the interpretation of the experimental findings. In this model, it is assumed that the only chemical species that passes through the X. laevis cell membrane is ${\rm CO}_{2}$, while all other chemical species that are present in the solution cannot cross the membrane as it is de facto impermeable to them. This hypothesis is supported by earlier studies, see, e.g., [16, 14].
@@
The standard experiment comprises the measurement of the surface pH on the cell membrane by a pH-sensitive electrode placed against the membrane. At the beginning of the experiment, the data show a steep rise of the surface pH caused by the association of protons and bicarbonate into carbonic acid that breaks into water and ${\rm CO}_{2}$ replacing the carbon dioxide passing through the membrane into the cell. As the experiment proceeds, the ${\rm CO}_{2}$ transport slows down as the concentration inside the cell rises towards that of the outside, and this, together with the diffusion from the exterior domain drives the system to a new equilibrium. The virtual proton transport into the cell is enhanced by the presence of carbonic anhydrases (CA), a family of enzymes that increase the dissociation of the ${\rm H}_{2}{\rm CO}_{3}$ right outside the cell membrane, while promoting the reverse reaction right under the membrane surface [5, 21, 24, 23], as illustrated schematically in Figure 1. The concentration and spatial distribution of CA vary for different kind of cells, and it is known to be significantly high, for example, in erythrocytes.
@@
To localize the effect of the membrane-bound carbonic anhydrase in the FEM model, it was necessary to resort to a dense spatial discretization, leading to a model with tens of thousands of degrees of freedom. Furthermore, because of the very fast dissociation of carbonic acid and the slow diffusion process, the characteristic time scales range over several orders of magnitude, making the semi-discretized dynamic system extremely stiff. Consequently, to capture the fast-slow dynamics, in [7] it was necessary to use an implicit-explicit scheme that separates different time scales. On a standard laptop or desktop computer, the resulting forward solver took typically several hours for each forward solve, making the algorithm impractical for estimating the membrane permeability and other parameters from the pH data. To overcome the issues of computational complexity, a reduced approximate model still accounting for the effect of the sensor was proposed in [2], and subsequently used for the estimation of the membrane permeability as the evolution step of a Bayesian particle filter algorithm in [3].
@@
In recent years, there has been a growing interest in using dictionary learning methods for solving inverse problems when the information in the data is not sufficient for the identification of any meaningful parametric model, or the forward model is not amenable for classical optimization-based approaches. Following this idea, here we propose a new alternative algorithm for solving the inverse problems of estimating the cell membrane properties based on a dictionary learning algorithm introduced in [4]. The main advantage of the new approach is that it is computationally much more efficient than the particle filter algorithm, estimating the desired parameters to a similar accuracy as the particle filter at a small fraction of the computing time.
@@ 2 Experimental framework and mathematical model
The experiment starts with a homogeneous solution containing carbon dioxide (${\rm CO}_{2}$), bicarbonate (${\rm HCO}_{3}^{-}$), carbonic acid (${\rm H}_{2}{\rm CO}_{3}$), hydrogen ion ($\rm H$$^{+}$), a non-specified pH buffer other than bicarbonate ($\rm A$$^{-}$), and the relative acid ($\rm HA$}) in equilibrium with known initial concentrations. In the standard experiment, the X. laevis oocyte is immersed in this solution, and the micro-electrode measuring the membrane surface pH is already in place. The interior of the cell is assumed to consist of a solution of the same species as the exterior, but with different known equilibrium concentrations.
@@
In [7], where the micro-environment under the tip of the sensor was modeled using a 3D finite element method approach, a substantial computational effort was required for each simulation, highlighting the need for a lower complexity model for the solution of the inverse problem of membrane permeability. Following [2], here we resort to a lumped compartment model for the perturbation caused by the sensor tip near the cell membrane. Based on the reasonable assumptions that the diffusion is mainly in the radial direction, and that the electrode has a minimal effect on the diffusion away from the membrane, the lumped model is effectively a radial diffusion model, ignoring for the most part diffusion in the direction parallel to the cell membrane, except right at the cell membrane, where diffusion from the free space to the micro-environment under the electrode tip is accounted for. The model is discussed in detail in the following paragraphs.
@@
where $u_{n}^{\pm}=u_{n}^{\pm}(x,t)$ is the concentration of the $n$th species at time $t$ and point $x$, the superscript $\pm$ indicates whether the concentration refers to that inside (-) or the outside (+) the cell membrane, $\varphi^{\pm}_{nj}=\varphi^{\pm}_{nj}(x,t)$ is the reaction flux of the $j$th reaction and, finally, $S_{nj}$ are the entries of the stoichiometric matrix $S$. More specifically, the numbering of the concentrations is $u_{1}=[{\rm CO}_{2}]$, $u_{2}=[{\rm H_{2}CO}_{3}]$, $u_{3}=[{\rm HCO}_{3}^{-}]$, $u_{4}=[{\rm H}^{+}]$, $u_{5}=[{\rm HA}]$, and $u_{6}=[{\rm A}^{-}]$. The diffusion coefficients $\kappa_{n}^{\pm}$, assumed to be known, are set to the values reported in the literature listed in Table 3.
@@
Substance Inside $(\mu\,{\rm m})^{2}/s$] Outside $(\mu\,{\rm m})^{2}/s$]
$\rm CO_{2}$} $\times 10^{3} 1.71$$\times 10^{3}$
$\rm H_{2}CO_{3}$} $\times 10^{3} 1.11$$\times 10^{3}$
$\rm HCO_{3}$$^{-}$ $\times 10^{3} 1.11$$\times 10^{3}$
$\rm H$$^{+}$ $\times 10^{3} 8.69$$\times 10^{3}$
$\rm HA$} $\times 10^{3} 1.56$$\times 10^{3}$
$\rm A$$^{-}$ $\times 10^{3} 1.56$$\times 10^{3}$
@@
At the beginning of the experiment, the systems inside and outside the membrane are in equilibrium with concentrations constants throughout the exterior domain and within the cell, however, the constant concentrations are not equal across the membrane boundary. The initial values of the concentrations are listed in Table 3. The assumption that the solution is well-mixed and the concentrations of the species far from the cell are well-known justifies assigning Dirichlet boundary conditions at an artificial outer boundary of the spherical computational domain with radius $R_{\infty}$,
@@
Reaction $_{\ell} kk_{-\ell}$ $K=k_{\ell}/k_{-\ell}$
[…]
${\rm H}_{2}{\rm CO}_{3}{\mathop{\rightleftharpoons}^{k_{2}}_{k_{-2}}}{\rm HCO}_{3}^{-}+{\rm H}^{+}$} $\varepsilon=10^{-9}$ [1/s] $\varepsilon/K_{2}$ $K_{2}=0.2407$ [mM]
${\rm HA}_{\rm in}{\mathop{\rightleftharpoons}^{k_{3}}_{k_{-{3}}}}{\rm A}^{-}_{\rm in}+{\rm H}^{+}_{\rm in}$} $\varepsilon^{\prime}=10^{-6}$ [1/s] $\varepsilon^{\prime}/K_{\rm HA}$ $K_{\rm HA}=7.9433\times 10^{-5}$ [mM]
${\rm HA}_{\rm out}{\mathop{\rightleftharpoons}^{k_{3}}_{k_{-{3}}}}{\rm A}^{-}_{\rm out}+{\rm H}^{+}_{\rm out}$} $\varepsilon^{\prime}=10^{-6}$ [1/s] $\varepsilon^{\prime}/K_{\rm HA}$ $K_{\rm HA}=3.1623\times 10^{-5}$ [mM]
Table 4: Reaction rates. Note that $\varepsilon$ and $\varepsilon^{\prime}$, whose precise values are not well known, represent the fast rates of the reactions for which no better guess is available. Furthermore, in the simulations we use $Ak_{\pm 1}$, as those reactions are enhanced by the (multiplicative) acceleration factor AA.
@@
$\gamma$ $=$ $\gamma_{0}\,10^{[-0.5\,\xi_{\gamma}+(1-\xi_{\gamma})]},\quad 0\leq\xi_{\gamma}\leq 1.$ (19)
@@ 3 Bayesian dictionary learning
In data science, dictionary learning and dictionary matching refer to a class of algorithms aiming at explaining a given datum in terms of a collection of precomputed or previously measured and labeled entries, referred to as the underlying dictionary. In applications where the new datum is a vector $b$ of measured response of a physical system, dictionary learning can be employed to explain it as a linear combination of a collection of possible responses. Assuming that the known responses are stored as vectors $d_{1},\dots,d_{p}$ of the same dimension as $b$, often referred to as atoms, they can be organized as the columns of a matrix dictionary $\mathsf{D}$}, and the problem can be formulated as finding a solution of the linear system
@@
In inverse problem applications, the atoms may represent precomputed outputs corresponding to a model that depends on model parameters that one is primarily interested in, the model parameter values serving as labels. Identifying one or few atoms that are able to explain the datum allows an identification of the model parameters, thus providing a solution to the inverse problem. The approach is particularly useful for problems in which the forward model is not amenable for classical methods, see, e.g., [20, 1] for examples. In the current application, the dictionary consists of precomputed pH curves, each corresponding to a given parameter value combination.
@@
For computational efficiency, it is often advantageous to preprocess the dictionary and identify possible clusters, i.e., divide the atoms in groups, each group consisting of atoms with similar characteristics. In lack of better criteria, one con resort to unsupervised clustering techniques such as $k$-means or $k$-medoids algorithm. Assuming that the dictionary atoms are partitioned into $K$ clusters, refereed to as sub-dictionaries, following [4], we can permute the columns of $\mathsf{D}$} so that columns corresponding to atoms in the same cluster are adjacent, yielding the block structure
@@
In this manner the dictionary learning problem can be solved by first identifying the sub-dictionaries that may be pertinent, and subsequently reformulating (20) with a smaller dictionary matrix that does not include the blocks corresponding to irrelevant sub-dictionaries. It has been shown in [4] that this approach can substantially reduce the computational complexity, especially when the dictionary comprises a large number of atoms.
In data science applications, it is customary to reduce the size of large data matrices by means of low-rank matrix approximations. If the columns of a data matrix have some degree of similarity, it is natural to identify a few vectors capturing their main characterizing features, and to approximate them as linear combinations of the feature vectors. This leads to a low rank representation of the data vectors. Once the sub-dictionaries have been identified, e.g., by clustering, each one can be compressed and approximated by just a few feature vectors. More specifically, if ${\mathsf{D}}^{(i)}\in{\mathbb{R}}^{m\times p_{i}}$, we write a low-rank approximation
@@
that is, every atom in the sub-dictionary ${\mathsf{D}}^{(i)}$ is approximated by a linear combination of the columns of $W^{(i)}$, referred to as the $i$th code book. The rank $k_{i}$ represents the effective dimensionality of the subdictionary, and $W^{(i)}$ represents a reduced sub-dictionary because it summarizes ${\mathsf{D}}^{(i)}$ in terms of much fewer feature vectors. Observe that clustering the the dictionary entries in subdictionaries of similar entries reduces the diversity within each group, thus helping to lower the number of feature vectors needed to explain each subdictionary. A popular low-rank approximation technique, the principal component analysis (PCA), is based on the singular value decomposition (SVD) of the matrix, however, if positivity constraints for the feature vectors and/or the coefficients are necessary, non-negative matrix factorization (NMF) algorithms may be preferable [19]. While for a given approximation rank $k_{i}$, the approximation properties of NMF measured in the Frobenius norm are inferior to those of PCA, the non-negativity of the feature vectors and of the weights facilitate the interpretation of the factorization.
@@
the inequalities being understood component-wise. The columns of the matrix ${\mathsf{W}}^{(i)}$, sometimes referred to as feature vectors, can be interpreted as a summary of the atoms of ${\mathsf{D}}^{(i)}$, thus ${\mathsf{W}}^{(i)}$ can be thought of as a compressed version of ${\mathsf{D}}^{(i)}$. In the current application, we compute the NMF using an iterated nonnegative least squares algorithm with sparsity constraints for the matrix ${\mathsf{H}}^{(i)}$ as discussed in more detail later.
@@
${\mathsf{D}}^{(i)}={\mathsf{W}}^{(i)}{\mathsf{H}}^{(i)}+{\mathsf{E}}^{(i)}\,$, (25)
where the discrepancy matrix ${\mathsf{E}}^{(i)}$ can be explicitly computed. In the proposed algorithm, we adopt a probabilistic view, thus the columns of the matrix ${\mathsf{D}}^{(i)}$, and therefore those of ${\mathsf{E}}^{(i)}$, are modeled as independent samples from some underlying probability distribution. In turn, if the columns of ${\mathsf{E}}^{(i)}$ are independent realizations of a random variable, denoted by $E^{(i)}$, we can estimate the underlying probability distribution from a sample of realizations of this random variable. Therefore, for each subdictionary, we consider the sample of dictionary entries $\{d_{j}^{(i)}\}_{j=1}^{K_{i}}$ comprising the columns of the matrix ${\mathsf{D}}^{(i)}$, find their approximate representations in terms of the feature vectors, and compute the approximation errors,
@@
$\widehat{i}=\underset{i}{\rm argmin}\left\|\left({\mathsf{C}}_{\rm DCE}^{(i)}\right)^{-1/2}\left(b-\mu^{(i)}_{\rm DCE}-{\mathsf{W}}^{(i)}h^{(i)}\right)\right\|\,.$ (30)
@@
The final step of the process entails going from the solution vector $\widehat{x}$ to the parameter vector $\xi$ associated with the measured data $b$. If $\widehat{x}$ contains only a single non-zero component, the parameter identification is straightforward. When the sparse vector $x$ contains more than one non-zero entry, a winner-takes-all strategy of electing the parameter vector corresponding to the best matching dictionary atom is a feasible approach. In the computed examples of this paper, we use an approximate interpolation scheme. Consider the mapping from the parameter space to the data space, $f:\xi\mapsto d$. We denote by $\xi_{j}^{(\hat{i})}$ the parameter vectors corresponding to the best matching subdictionary, $1\leq j\leq p_{\,\hat{i}}$, and by $d_{j}^{(\hat{i})}$ the corresponding computed outputs. We normalize the coefficients $\widehat{x}_{j}$, defining
$\eta_{j}=\frac{1}{\|\hat{x}\|_{1}}\hat{x}_{j},$
and define the convex combination of the parameter vectors,
$\overline{\xi}=\sum_{j=1}^{p_{\hat{i}}}\eta_{j}\xi_{j}^{(\hat{i})}$
Assuming that the individual parameter vectors $\xi^{(\hat{i})}_{j}$ are close to the convex combination $\overline{\xi}$, we obtain by linearization around $\overline{\xi}$ the approximate identity
[…]
$\approx$ $\frac{1}{\|\hat{x}\|_{1}}\sum_{j=1}^{p_{\,\widehat{i}}}\eta_{j}\left(f(\overline{\xi})+Df(\overline{\xi})(\xi^{(\hat{i})}_{j}-\overline{\xi})\right)$
$=$ $\frac{1}{\|\widehat{x}\|_{1}}\left(f(\overline{\xi})+Df(\overline{\xi})\big(\sum_{j=1}^{p_{\,\hat{i}}}\eta_{j}\xi^{(\hat{i})}_{j}-\overline{\xi}\big)\right)$
$=$ $\frac{1}{\|\widehat{x}\|_{1}}f(\overline{\xi}).$
This calculation indicates that the best matching parameter value explaining the data is given by the convex combination
$\widehat{\xi}=\sum_{j=1}^{p_{\,\hat{i}}}\eta_{j}\xi_{j}^{(\hat{i})}$. (32)